# Communications, Computer engineering, New computational paradigms (quantum)

### v2

@@ Abstract
The Quantum Internet is key for distributed quantum computing, by interconnecting multiple quantum processors into a virtual quantum computation system. This allows to scale the number of qubits, by overcoming the inherent limitations of noisy-intermediate-scale quantum (NISQ) devices. Thus, the Quantum Internet is the foundation for large-scale, fault-tolerant quantum computation. Among the distributed architectures, Quantum Data Centers emerge as the most viable in the medium-term, since they integrate multiple quantum processors within a localized network infrastructure, by allowing modular design of quantum networking. We analyze the physical and topological constraints of Quantum Data Centers, by emphasizing the role of entanglement orchestrators in dynamically reconfiguring network topologies through local operations. We examine the major hardware challenge of quantum transduction, essential for interfacing heterogeneous quantum systems. Furthermore, we explore how interconnecting multiple Quantum Data Centers could enable large-scale quantum networks. We discuss the topological constraints of such a scaling and identify open challenges, including entanglement routing and synchronization. The carried analysis positions Quantum Data Centers as both a practical implementation platform and strategic framework for the future Quantum Internet.
keywords:
Quantum Data Center, entanglement, quantum network
@@ 1 Introduction
Quantum computing is one of the most promising applications of quantum technologies, gathering significant investments from leading technology firms such as Google, IBM, and Amazon. Indeed, the promises of quantum computing span across all major industrial sectors, with particular emphasis on sustainable energy optimization, an area where quantum-enabled advances are expected to drive significant progress in the coming years. Furthermore, as Shor demonstrated in 1994 [Shor1994], quantum computing has the potential to break Rivest–Shamir–Adleman (RSA), the most widely cryptographic protocol used on Internet nowadays. And for its implementation, thousands of qubits are required [CacCalTaf-20, CalAmoFer-22]. More into details, breaking a $2048$-bit RSA key requires roughly $4\text{,}099$ logical (or fault-tolerant) [Bea-03] qubits. The mapping between logical qubits and physical qubits depends on the specific adopted technology (or qubit implementation) and usually implies a huge overhead. By considering the different qubit implementations and current projections, it is estimated that $10^{4}$ to $10^{7}$ interconnected physical qubits are required to break current cryptography standards [mckinsey-24]. However, in the state-of-the-art quantum computers, the number of physical qubits is limited to double digits due to interference, cross-talking, over-heating and challenges in quantum control systems [Pre-18].
To overcome the aforementioned scalability challenges, multiple quantum processors – with a limited number of physical qubits – can be interconnected through a quantum communication network, by adopting a distributed computing approach. This strategy mirrors the classical distributed computing model, where a computation is properly partitioned so that multiple processors can cooperate over a network to solve complex tasks that choke single machine, by leveraging shared data and resources. Surprisingly, unlike classical systems – where the computational power scales linearly with the number of interconnected computing modules – in the quantum domain the computational power can scale exponentially [CacCalTaf-20], thanks to entanglement. For this reason, both the research community and leading industry players are currently focusing on the development of quantum networks for enabling distributed quantum computing. Notably, IBM has already successfully linked two quantum processors and announced plans to scale this to a seven-QPU network.[CarVaz2024].
@@ 1.1 Entanglement-enabled communications
In distributed quantum computing, qubits are physically separated across multiple processors. Thus, when a quantum gate must act on remote qubits, namely on non-co-located qubits, specialized communication primitives are required to enable inter-processor operations [CalAmoFer-22]. However, these primitives cannot be borrowed from classical distributed computing model, since the underlying quantum physics demands a fundamentally different paradigm.
@@
Unlike classical bits, qubits are subjected to quantum decoherence, a phenomenon that irreversibly scrambles quantum states and therefore the encoded information [CalAmoFer-22]. This kind of quantum noise affects every stage of the quantum distributed computing: processing, storage, and transmission. Accordingly qubits cannot be stored indefinitely, but must be processed within the coherence time. Furthermore, differently from classical distributed paradigm, the direct transfer of quantum data from one processor to another is not easy, due to the no-cloning theorem, which forbids replicating unknown quantum states. As a consequence, the direct transmission of a qubit between processors is inherently risky. Indeed, if a qubit carrying quantum information is lost or corrupted during transmission on the physical channel, its quantum state – and thus the encoded information– is irreversibly lost. This makes direct transmission unreliable.
Thankfully, quantum entanglement, a unique quantum phenomenon where particles share correlated states regardless of the distance, can be exploited as the key communication resource to avoid the issues arising with the direct transmission of qubits. More into details, entangled states, such as Bell pairs, can be distributed in advance through a physical quantum channel [^1] Since entangled states are communication resources and do not encode information, they are known states. Thus, they are not subjected to the limitations of the no-cloning theorem. As a result, repeated transmission attempts are allowed, until successful distribution is achieved.. Once established, the entangled pair enables the quantum teleportation protocol [CacCalVan-20], which leverages the non-local effects activated by the shared entanglement to “transmit” the informational qubit without the physical transfer of the particle encoding information. These principles form the foundation of the computational paradigms known as TeleData and TeleGate, which respectively allow the transfer of quantum data and quantum gates across distributed quantum systems [CalAmoFer-22]. These primitives generalize the concept of moving quantum states across multiple processors in distributed quantum computing, by avoiding the unreliability of direct quantum information transmission. In a nutshell, entangled states, once shared, serve as a virtual link, enabling remote operations to be performed across multiple non-co-located qubits, as they resided within the same quantum processor.
@@
Figure 1: High-level representation of the Distributed Quantum Computing. A crucial in-the-middle component is the compiler, an entity responsible of translating and hardware-agnostic description of the algorithm into a partitioned and fitted version to be executed over the set of individual quantum processors [CuoCalKrs-21, FerCacAmo-21]. As represented in the lowest part of the figure, the set of quantum processing units act as a virtual quantum processor thanks to the interplay of several entities comprising the network infrastructure, represented as quantum communication channels and classical channels, quantum processors and classical processors. Remarkably, the distributed landscape is enabled though three main entanglement-based functional blocks: entanglement generation, entanglement distribution and entanglement utilization.
The remaining part of this work is organized as follows. In Sec. 2 describes the infrastructural archetypes of quantum networks designed for the distributed quantum computing with a focus on Quantum Data Centers. Then, in Sec. 3, we provide a detailed discussion on the operational principles, the physical constraints, and topological challenges of Quantum Data Centers from a communication engineering perspective. In Sec. 4 future perspectives on the distributed quantum computing landscape are presented with a discussion on the road beyond Quantum Data Centers. Specifically, we discuss the communication requirements for the development of the so-called Quantum Hubs. Finally, Sec. 5 concludes the paper with open issues and further insights.
@@ 2 Distributed quantum computing: Archetypes
By enriching with details the abstraction represented in Fig. 1, three main archetypes for distributed quantum computing can be recognized [CalAmoFer-22]: Multi-Core quantum computer, Quantum Data Centers and Quantum Hub. As summarized in Fig. 2, the three archetypes differ each others for three main features: the scale of involved processing entities, the complexity of the network infrastructure, enabling the distributed computation and the technology heterogeneity.
Multi-core archetype is characterized by the lowest values across the feature indicators. As detailed in [CalAmoFer-22, RusVinPal-25], this archetype marks the boundary between single-core quantum computing architectures and distributed quantum computing architectures. In the Multi-Core archetype, multiple quantum processing units are interconnected within a single quantum computer. The most straightforward example of Multi-Core architecture is the deployment of multiple superconducting-based quantum chips within the same cryostat or through advanced couplers [ManCorGam-24, Abu-25]. This archetype exhibits the potential of increasing the number of interconnected physical qubit at a relatively low-cost and efforts, being characterized by low-to-no heterogeneity. Furthermore, Multi-Core architectures exhibit low-network complexity as the network connections can be hard-designed, by leveraging chip-scale interconnections. This, in turn, implies that the number of interconnected cores is inevitably bounded by the rack/refrigerator volume. Thus, the number of physical qubits that can be clustered together is limited as well. Accordingly, the network functionalities are also relatively simple[CalAmoFer-22, EscRacRod-23].
As represented in Fig. 2, the Quantum Hub, namely the distributed quantum computing configuration interconnecting multiple diverse Quantum Data Centers, is the most complex archetype in our classification, being characterized by the highest grade of network complexity, heterogeneity and scale of the interconnected entities. Indeed, significant heterogeneity may arise, since different Quantum Data Centers are likely owned by different operators or rely on different hardware platforms. Furthermore, the interconnection of geographically-distributed centers requires a wide-scale network infrastructure, likely enabled by the Quantum Internet [CalAmoFer-22, Kim-08, rfc9340]. In this context, it is clear that network interoperability across platforms should be supported, while ensuring a reliable management of entanglement generation and distribution at large scale. This, in turn, requires seamless integration with the classical Internet infrastructure [CacIllKou-22, CacIllCal-23].
@@
2.1 Multi-Computer Archetype: Quantum Data Center
Quantum Data Centers represent the bridge between the small-scale distributed quantum computing systems and the large-scale, high-complexity archetype [LiuJia-24, LiuHanJia-23-1, AghAinAle-25]. In this intermediate configuration, multiple quantum computers – each comprising of multiple quantum cores – are located in the same controlled environment (e.g. a dedicated facility) and interconnected via a Quantum Local Area Network (QLAN) [MazCalCac-25, MazCalCac-24-QCNC, MazZhaChu-24].
@@ 2.1 Multi-Computer Archetype: Quantum Data Center
As aforementioned in the introduction, the design and development of Quantum Data Centers represent the next quantum engineering challenge, as already acknowledged by the major stakeholders and tech players in the related areas[Cisco-24, ManCorGam-24, ibm2024]. For this, the remaining part of the paper focuses on the architectural and network operational challenges of Quantum Data Centers from a communication engineering perspective. Before concluding this section, it is worthwhile to mention that given the inherent complexities of the Quantum Data Center archetype, we envision both classical and quantum Machine Learning (ML) techniques to play a relevant role in its evolution. Indeed classical and quantum ML techniques can be exploited for managing and optimizing critical processes, including quantum control, error mitigation, and entanglement distillation strategies[NeuWez-22, BenLloSac-19, ChiSim-23, PhiChi-22, ChiHar-23, AdeKum-25, LocCarMul-23].
3 Quantum Data Centers: constraints on implementation and physical topology
As discussed, entanglement serves as fundamental resource enabling quantum operations between qubits distributed across different quantum computers within a Quantum Data Center. However, its implications extend far beyond this functionality. Indeed, entanglement revolutionizes the very same concept of a communication network, empowering an entirely new concept of connectivity, transcending the classical networking limitations[IllCalMan-22, CalAmoFer-22, MazCalCac-25, MazZhaChu-24, CheIllCac-24]. Consequently, the ability to decouple the physical QLAN network topology from the actual capacity to exchange quantum information in a Quantum Data Center, via the entanglement-activated connectivity, introduces novel and largely unexplored network design paradigms[IllCalVis-23, CalAmoFer-22, MazZhaChu-24, MazCalCac-25]. More into details, QLAN deployments require resource-intensive, costly, and sophisticated setups [DavCacCal-2024, CalDavHan-25, MazCalCac-25]. This, in turn, implies, at least in the mid-term time horizon, that:
  • the QLAN physical topology significantly diverges from the highly connected, densely populated, and highly dynamic classical data center topologies [MazCalCac-25];
  • the ability to generate, distribute, and manipulate entangled states is confined into specialized nodes. In fact, technology maturity prevents homogeneous node functionalities[RuiWalDur-25, DavCacCal-2024, DavZhaChu-24].
These constraints necessitate a hierarchical QLAN design where specialized nodes, referred to as the orchestrator nodes, are responsible for the entanglement generation and distribution [DavCacCal-24, CheIllCac-24, MazCalCac-25], according to different strategies [CacIllVis-24]. Thus, as extensively discussed later in this section, in the orchestrator node is concentrated the complexity, by handling the most resource-intensive network operations. While the remaining network nodes are simple, by exhibiting limited capabilities.
@@ 3.1 Quantum Transduction
Quantum Data Centers inherently rely on heterogeneous qubit platforms, since no single qubit technology currently meets all the requirements needed for quantum computation and communication, as detailed in the following.
Superconducting technology stands out as a leading platform for quantum computations for its key advantages. In fact, fast, high-fidelity gates can be realized via this hardware platform, by ensuring high circuit scalability. Furthermore, the superconducting technology is characterized by efficient control and readout mechanisms [KjaSchBra-20, Wen-17, OfePetRei-16]. However, these benefits come at the price of superconducting qubits operating at cryogenic temperatures (few tens of mK), which makes them very susceptible to thermal noise and decoherence.
With this in mind, for connecting superconducting nodes in a Quantum Data Center for distributed quantum computation, in principle cryogenic waveguides [XiaZhaLia-17, KurPechRoy-19, ManStoKur-20] can be utilized. However, cryogenic cables present significant practical limitations. Indeed, they are characterized by very expensive manufacturing costs and non-trivial installation requirements. Furthermore, they introduce severe communication constraints, such as very limited communication range and inflexible network topology. This, in turn, results in huge difficulties in reconfiguring the network architecture. Consequently, purely cryogenic interconnection strategies are not viable for scaling distributed quantum computing systems.
@@
An alternative solution for interconnecting superconducting processors leverages an optical fiber network for entanglement distribution. Indeed, quantum optics technology offers several advantages: weak interaction with the environment (low decoherence), high-fidelity transmission, easy control with standard optical components and high-speed operations [RenXuYon-17, CacCalTaf-20]. Therefore, optical photons, also referred to as flying qubits, are widely recognized as entanglement carriers.
Compared to cryogenic-based interconnections, fiber-based solutions offer key several assets not limited to the significantly lower infrastructure costs. Specifically, they enable long-range communications, and accommodate diverse network topologies, by allowing dynamic reconfigurations of the network architecture. In addition, a compelling advantage of optical quantum networks is their ability to leverage existing telcom-fiber infrastructures. In fact, recent results show that entanglement distribution can be achieved by utilizing lit fibers, without additional costs, by eliminating the need of dedicated quantum channels, such as dark fibers. Clearly, for a seamless integration with classical networks, the coexistence of classical and quantum signals – sharing the same fiber – has to maintain quantum integrity via noise suppression techniques [JorFeiChe-24]. The above comparison is summarized [^2] We kindly refer the interested reader to [CalDavHan-25] for a thorough and comprehensive discussion on quantum transduction from a communication perspective in Tab.1.
However, the crucial obstacle for interconnecting superconducting nodes with an optical QLAN network stems from the huge frequency gap between optical photons (that typically work at about hundred of THz) and superconducting qubits (that work at a few GHz). This five-order-of-magnitude gap needs Quantum Transduction [CacCalTaf-20, LauSinBar-20], i.e. the process of converting one type of qubit to another, by enabling the interaction of different qubit platforms [CalDavHan-25]. A quantum transducer, namely the network component performing quantum transduction, constitutes a mandatory interface converting a superconducting qubit within a network node into a flying qubit that travels through optical channels and vice versa [CalDavHan-25, DavCacCal-24], as schematically depicted in Fig. 3. However, as clarified in the next paragraph, enabling an effective quantum transduction, by preserving the quantum states, is still an open problem for the current technology maturity. Therefore, quantum transduction remains the main limitation in utilizing optical fibers for interconnecting superconducting nodes in a Quantum Data Center.

### v3

# Quantum Data Centres:

@@ Abstract
The Quantum Internet is key for distributed quantum computing, by interconnecting multiple quantum processors into a virtual quantum computation system. This allows to scale the number of qubits, by overcoming the inherent limitations of noisy-intermediate-scale quantum (NISQ) devices. Thus, the Quantum Internet is the foundation for large-scale, fault-tolerant quantum computation. Among the distributed architectures, Quantum Data Centres emerge as the most viable in the medium-term, since they integrate multiple quantum processors within a localized network infrastructure, by allowing modular design of quantum networking. We analyze the physical and topological constraints of Quantum Data Centres, by emphasizing the role of entanglement orchestrators in dynamically reconfiguring network topologies through local operations. We examine the major hardware challenge of quantum transduction, essential for interfacing heterogeneous quantum systems. Furthermore, we explore how interconnecting multiple Quantum Data Centres could enable large-scale quantum networks. We discuss the topological constraints of such a scaling and identify open challenges, including entanglement routing and synchronization. The carried analysis positions Quantum Data Centres as both a practical implementation platform and strategic framework for the future Quantum Internet.
Keywords Quantum Data Centre, entanglement, quantum network, QNattyNet
@@ 1 Introduction
Quantum computing is one of the most promising applications of quantum technologies, gathering significant investments from leading technology firms such as Google, IBM, and Amazon. Indeed, the promises of quantum computing span across all major industrial sectors, with particular emphasis on sustainable energy optimization, an area where quantum-enabled advances are expected to drive significant progress in the coming years. Furthermore, as Shor demonstrated in 1994 [66], quantum computing has the potential to break Rivest–Shamir–Adleman (RSA), the most widely cryptographic protocol used on Internet nowadays. And for its implementation, thousands of qubits are required [10, 15]. More into details, breaking a $2048$-bit RSA key requires roughly $4\text{,}099$ logical (or fault-tolerant) [7] qubits. The mapping between logical qubits and physical qubits depends on the specific adopted technology (or qubit implementation) and usually implies a huge overhead. By considering the different qubit implementations and current projections, it is estimated that $10^{4}$ to $10^{7}$ interconnected physical qubits are required to break current cryptography standards [23]. However, in the state-of-the-art quantum computers, the number of physical qubits is limited to double digits due to interference, cross-talking, over-heating and challenges in quantum control systems [60].
To overcome the aforementioned scalability challenges, multiple quantum processors – with a limited number of physical qubits – can be interconnected through a quantum communication network, by adopting a distributed computing approach. This strategy mirrors the classical distributed computing model, where a computation is properly partitioned so that multiple processors can cooperate over a network to solve complex tasks that choke single machine, by leveraging shared data and resources. Surprisingly, unlike classical systems – where the computational power scales linearly with the number of interconnected computing modules – in the quantum domain the computational power can scale exponentially [10], thanks to entanglement. For this reason, both the research community and leading industry players are currently focusing on the development of quantum networks for enabling distributed quantum computing. Notably, IBM has already successfully linked two quantum processors and announced plans to scale this to a seven-QPU (Quantum Processing Unit) network.[17].
@@ 1.1 Entanglement-enabled communications
In distributed quantum computing, qubits are physically separated across multiple processors. Thus, when a quantum gate must act on remote qubits, namely on non-co-located qubits, specialized communication primitives are required to enable inter-processor operations [15]. However, these primitives cannot be borrowed from classical distributed computing model, since the underlying quantum physics demands a fundamentally different paradigm.
@@
Unlike classical bits, qubits are subjected to quantum decoherence, a phenomenon that irreversibly scrambles quantum states and therefore the encoded information [15]. This kind of quantum noise affects every stage of the quantum distributed computing: processing, storage, and transmission. Accordingly qubits cannot be stored indefinitely, but must be processed within the coherence time. Furthermore, differently from classical distributed paradigm, the direct transfer of quantum data from one processor to another is not easy, due to the no-cloning theorem, which forbids replicating unknown quantum states. As a consequence, the direct transmission of a qubit between processors is inherently risky. Indeed, if a qubit carrying quantum information is lost or corrupted during transmission on the physical channel, its quantum state – and thus the encoded information– is irreversibly lost. This makes direct transmission unreliable.
Thankfully, quantum entanglement, a unique quantum phenomenon where particles share correlated states regardless of the distance, can be exploited as the key communication resource to avoid the issues arising with the direct transmission of qubits. More into details, entangled states, such as Bell pairs, can be distributed in advance through a physical quantum channel [^1] Since entangled states are communication resources and do not encode information, they are known states. Thus, they are not subjected to the limitations of the no-cloning theorem. As a result, repeated transmission attempts are allowed, until successful distribution is achieved.. Once established, the entangled pair enables the quantum teleportation protocol [11], which leverages the non-local effects activated by the shared entanglement to “transmit” the informational qubit without the physical transfer of the particle encoding information. These principles form the foundation of the computational paradigms known as TeleData and TeleGate, which respectively allow the transfer of quantum data and quantum gates across distributed quantum systems [15]. These primitives generalize the concept of moving quantum states across multiple processors in distributed quantum computing, by avoiding the unreliability of direct quantum information transmission. In a nutshell, entangled states, once shared, serve as a virtual link, enabling remote operations to be performed across multiple non-co-located qubits, as they resided within the same quantum processor.
@@
Figure 1: High-level representation of the Distributed Quantum Computing. A crucial in-the-middle component is the compiler, an entity responsible of translating and hardware-agnostic description of the algorithm into a partitioned and fitted version to be executed over the set of individual quantum processors [24, 31]. As represented in the lowest part of the figure, the set of quantum processing units act as a virtual quantum processor thanks to the interplay of several entities comprising the network infrastructure, represented as quantum communication channels and classical channels, quantum processors and classical processors. Remarkably, the distributed landscape is enabled though three main entanglement-based functional blocks: entanglement generation, entanglement distribution and entanglement utilization.
The remaining part of this work is organized as follows. In Sec. 2 describes the infrastructural archetypes of quantum networks designed for the distributed quantum computing with a focus on Quantum Data Centres. Then, in Sec. 3, we provide a detailed discussion on the operational principles, the physical constraints, and topological challenges of Quantum Data Centres from a communication engineering perspective. In Sec. 4 future perspectives on the distributed quantum computing landscape are presented with a discussion on the road beyond Quantum Data Centres. Specifically, we discuss the communication requirements for the development of the so-called Quantum Hubs. Finally, Sec. 5 concludes the paper with open issues and further insights.
@@ 2 Distributed quantum computing: Archetypes
By enriching with details the abstraction represented in Fig. 1, three main archetypes for distributed quantum computing can be recognized [15]: Multi-Core quantum computer, Quantum Data Centres and Quantum Hub. As summarized in Fig. 2, the three archetypes differ each others for three main features: the scale of involved processing entities, the complexity of the network infrastructure, enabling the distributed computation and the technology heterogeneity.
Multi-core archetype is characterized by the lowest values across the feature indicators. As detailed in [15, 63], this archetype marks the boundary between single-core quantum computing architectures and distributed quantum computing architectures. In the Multi-Core archetype, multiple quantum processing units are interconnected within a single quantum computer. The most straightforward example of Multi-Core architecture is the deployment of multiple superconducting-based quantum chips within the same cryostat or through advanced couplers [48, 1]. This archetype exhibits the potential of increasing the number of interconnected physical qubit at a relatively low-cost and efforts, being characterized by low-to-no heterogeneity. Furthermore, Multi-Core architectures exhibit low-network complexity as the network connections can be hard-designed, by leveraging chip-scale interconnections. This, in turn, implies that the number of interconnected cores is inevitably bounded by the rack/refrigerator volume. Thus, the number of physical qubits that can be clustered together is limited as well. Accordingly, the network functionalities are also relatively simple[15, 30].
As represented in Fig. 2, the Quantum Hub, namely the distributed quantum computing configuration interconnecting multiple diverse Quantum Data Centres, is the most complex archetype in our classification, being characterized by the highest grade of network complexity, heterogeneity and scale of the interconnected entities. Indeed, significant heterogeneity may arise, since different Quantum Data Centres are likely owned by different operators or rely on different hardware platforms. Furthermore, the interconnection of geographically-distributed centres requires a wide-scale network infrastructure, likely enabled by the Quantum Internet [15, 38, 40]. In this context, it is clear that network interoperability across platforms should be supported, while ensuring a reliable management of entanglement generation and distribution at large scale. This, in turn, requires seamless integration with the classical Internet infrastructure [13, 12].
@@
2.1 Multi-Computer Archetype: Quantum Data Centre
Quantum Data Centres represent the bridge between the small-scale distributed quantum computing systems and the large-scale, high-complexity archetype [45, 44, 3]. In this intermediate configuration, multiple quantum computers – each comprising of multiple quantum cores – are located in the same controlled environment (e.g. a dedicated facility) and interconnected via a Quantum Local Area Network (QLAN) [50, 49, 51].
@@ 2.1 Multi-Computer Archetype: Quantum Data Centre
As aforementioned in the introduction, the design and development of Quantum Data Centres represent the next quantum engineering challenge, as already acknowledged by the major stakeholders and tech players in the related areas[58, 48, 61]. For this, the remaining part of the paper focuses on the architectural and network operational challenges of Quantum Data Centres from a communication engineering perspective. Before concluding this section, it is worthwhile to mention that given the inherent complexities of the Quantum Data Centre archetype, we envision both classical and quantum Machine Learning (ML) techniques to play a relevant role in its evolution. Indeed classical and quantum ML techniques can be exploited for managing and optimizing critical processes, including quantum control, error mitigation, and entanglement distillation strategies[56, 8, 22, 59, 21, 2, 46].
3 Quantum Data Centres: constraints on implementation and physical topology
As discussed, entanglement serves as fundamental resource enabling quantum operations between qubits distributed across different quantum computers within a Quantum Data Centre. However, its implications extend far beyond this functionality. Indeed, entanglement revolutionizes the very same concept of a communication network, empowering an entirely new concept of connectivity, transcending the classical networking limitations[36, 15, 50, 51, 20]. Consequently, the ability to decouple the physical QLAN network topology from the actual capacity to exchange quantum information in a Quantum Data Centre, via the entanglement-activated connectivity, introduces novel and largely unexplored network design paradigms[37, 15, 51, 50]. More into details, QLAN deployments require resource-intensive, costly, and sophisticated setups [25, 16, 50]. This, in turn, implies, at least in the mid-term time horizon, that:
  • the QLAN physical topology significantly diverges from the highly connected, densely populated, and highly dynamic classical data centre topologies [50];
  • the ability to generate, distribute, and manipulate entangled states is confined into specialized nodes. In fact, technology maturity prevents homogeneous node functionalities[55, 25, 28].
These constraints necessitate a hierarchical QLAN design where specialized nodes, referred to as the orchestrator nodes, are responsible for the entanglement generation and distribution [26, 20, 50], according to different strategies [14]. Thus, as extensively discussed later in this section, in the orchestrator node is concentrated the complexity, by handling the most resource-intensive network operations. While the remaining network nodes are simple, by exhibiting limited capabilities.
@@ 3.1 Quantum Transduction
Quantum Data Centres inherently rely on heterogeneous qubit platforms, since no single qubit technology currently meets all the requirements needed for quantum computation and communication, as detailed in the following.
Superconducting technology stands out as a leading platform for quantum computations for its key advantages. In fact, fast, high-fidelity gates can be realized via this hardware platform, by ensuring high circuit scalability. Furthermore, the superconducting technology is characterized by efficient control and readout mechanisms [39, 69, 57]. However, these benefits come at the price of superconducting qubits operating at cryogenic temperatures (few tens of mK), which makes them very susceptible to thermal noise and decoherence.
With this in mind, for connecting superconducting nodes in a Quantum Data Centre for distributed quantum computation, in principle cryogenic waveguides [70, 41, 47] can be utilized. However, cryogenic cables present significant practical limitations. Indeed, they are characterized by very expensive manufacturing costs and non-trivial installation requirements. Furthermore, they introduce severe communication constraints, such as very limited communication range and inflexible network topology. This, in turn, results in huge difficulties in reconfiguring the network architecture. Consequently, purely cryogenic interconnection strategies are not viable for scaling distributed quantum computing systems.
@@
An alternative solution for interconnecting superconducting processors leverages an optical fiber network for entanglement distribution. Indeed, quantum optics technology offers several advantages: weak interaction with the environment (low decoherence), high-fidelity transmission, easy control with standard optical components and high-speed operations [62, 10]. Therefore, optical photons, also referred to as flying qubits, are widely recognized as entanglement carriers.
Compared to cryogenic-based interconnections, fiber-based solutions offer key several assets not limited to the significantly lower infrastructure costs. Specifically, they enable long-range communications, and accommodate diverse network topologies, by allowing dynamic reconfigurations of the network architecture. In addition, a compelling advantage of optical quantum networks is their ability to leverage existing telcom-fiber infrastructures. In fact, recent results show that entanglement distribution can be achieved by utilizing lit fibers, without additional costs, by eliminating the need of dedicated quantum channels, such as dark fibers. Clearly, for a seamless integration with classical networks, the coexistence of classical and quantum signals – sharing the same fiber – has to maintain quantum integrity via noise suppression techniques [67]. The above comparison is summarized [^2] We kindly refer the interested reader to [16] for a thorough and comprehensive discussion on quantum transduction from a communication perspective in Tab.1.
However, the crucial obstacle for interconnecting superconducting nodes with an optical QLAN network stems from the huge frequency gap between optical photons (that typically work at about hundred of THz) and superconducting qubits (that work at a few GHz). This five-order-of-magnitude gap needs Quantum Transduction [10, 42], i.e. the process of converting one type of qubit to another, by enabling the interaction of different qubit platforms [16]. A quantum transducer, namely the network component performing quantum transduction, constitutes a mandatory interface converting a superconducting qubit within a network node into a flying qubit that travels through optical channels and vice versa [16, 26], as schematically depicted in Fig. 3. However, as clarified in the next paragraph, enabling an effective quantum transduction, by preserving the quantum states, is still an open problem for the current technology maturity. Therefore, quantum transduction remains the main limitation in utilizing optical fibers for interconnecting superconducting nodes in a Quantum Data Centre.